OpenStudy (firejay5):

Again this is finding Holes & Vertical Asymptotes. Check my answer to see if they are correct to the following problems: *Medal Rewarded* 1. 4 over x^2 + 3x - 10 2. 2x^2 - x - 10 over 2x - 5 3. x^2 - x - 12 over x^2 - 4x 4. 3x - 1 over 3x^2 + 5x - 2 5. x^2 - 6x - 7 over x^2 + 6x - 7 6. 3x^2 - 5x - 2 over x + 3 7. x + 1 over x^2 - 6x + 5 8. 2x^2 - x - 3 over 2x^2 + 3x - 9

OpenStudy (firejay5):

My Answers: Some answers may be wrong or confusing 1. VA: x = 2 & x = -5 2. Hole: x = 5/2 3. VA: x = 0 & Hole: x = 4 4. VA: x = -2 & Hole: x = 1/3 5. VA: x = 1 & x = -7 6. VA: x = -3 7. VA: x = 5 & x = 1 8. VA: x = -3 & Hole: x = 2/3

jimthompson5910 (jim_thompson5910):

problems #1 through #4 are correct

jimthompson5910 (jim_thompson5910):

still checking the rest

OpenStudy (firejay5):

ok

jimthompson5910 (jim_thompson5910):

5 through 7 are correct

jimthompson5910 (jim_thompson5910):

8 is incorrect

OpenStudy (firejay5):

all of them are correct?

jimthompson5910 (jim_thompson5910):

everything but 8

OpenStudy (firejay5):

It's 3/2, NOT 2/3

jimthompson5910 (jim_thompson5910):

good catch

OpenStudy (firejay5):

I had it right, but it switched numbers

jimthompson5910 (jim_thompson5910):

once you fix that, everything else is perfect

OpenStudy (firejay5):

Are you sure 100%???

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (firejay5):

thanks so much! :D

jimthompson5910 (jim_thompson5910):

np

OpenStudy (firejay5):

a medal for that

jimthompson5910 (jim_thompson5910):

thx

OpenStudy (firejay5):

you gave me a medal lol! :D